waths the solution to 4x - 2x=6

Found 2 solutions by CertifiedGenius, superkid208:
Answer by CertifiedGenius(6)

What is the solution to 4x - 2x = 6?
I assume we are solving for x.
4x - 2x = 6 [given]
2x = 6 [subtract 2x from 4x to get 2x]
x = 3 [divide by 2 on each side to solve for x]
Now we can check our answer. Plug in 3 for every x.
(4*3) - (2*3) = 6 [given with substitutions for x]
12 - 6 = 6 [multiply 4 time 3 for 12 and 2 times 3 for 6]
6 = 6 [subtract 6 from 12 to make 6]

Answer by superkid208(3)

4x-2x=6
4x-2x=2x
so 2x=6
what times 2 = 6
so 2*3= 6
so x=3
4*3 - 2*3 =6
4*3=12 -2*3=6 =6
12-6=6
so x=3